OpenStudy (anonymous):

Tickets for a football match are sold at $30 for EACH adult and $15 for EACH child. A company bought 28 tickets. If x of these tickets were for adults, write an expression in terms of x for: a) the number of tickets for children b) the amount spent on tickets for adults c) the amount spent on tickets for children ii) Show that the TOTAL amount spent on the 28 tickets is $(15x +420) iii) The cost of the 28 tickets was $660, calculate the number of adult tickets bought by the company

OpenStudy (lgbasallote):

go @ParthKohli and @Arnab09 :DDD

OpenStudy (lgbasallote):

let's see who gives the better explanation >:))

OpenStudy (anonymous):

there are total 28 tickets. x are for adults, rest r for children so, there are 28-x tickets for children expense for children: multiply the number of children tickets with their cost, ie $15 so, expense for children: $15(28-x) expense for adults= x*price for each adult ticket= $30x so, total expense= $15(28-x)+$30x=$(15x+420)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (a) 28 -x}\) \(\Large \color{purple}{\rightarrow (b) 30x }\) \(\Large \color{purple}{\rightarrow (c) 15(28 -x ) }\) \(\Large \color{purple}{\rightarrow (i) 30x + 15(28 - x) = 30x + 420 - 15x = 15x + 420 }\) \(\Large \color{purple}{\rightarrow (ii) 15x + 420 = 660 }\) \(\Large \color{purple}{\rightarrow 15x = 240}\) \(\Large \color{purple}{\rightarrow x = 16 }\) So, 16 adult tickets. If you have any doubt, please ask.

OpenStudy (anonymous):

for the last one, total expense= 15x+420 it is given, total expense=$660 so, 15x+420=660 x=(660-420)/15=16 there are 16 adults, 28-26=12 children

Parth (parthkohli):

@lgbasallote Please point any mistake if there :p

OpenStudy (anonymous):

no, mistake :D

OpenStudy (lgbasallote):

i dont know how you got your answers parth =_= i said explanation :P

OpenStudy (anonymous):

eliminate the comma, BTW ^^

Parth (parthkohli):

Okay, first one is obvious. If there are some tickets bought by adults and total are 28, then 28 - x would be for children.

Parth (parthkohli):

2nd one, see, we have to multiply the price by the number of tickets, so it's 30 * x = 30x

Parth (parthkohli):

The 3rd one is almost the same, price * number of tickets.

Parth (parthkohli):

Total amount spent is amount for children + amount of adults, which I've written and solved.

Parth (parthkohli):

Last one: We know that the total amount spent is 660 dollars, so we know that 15x + 420 is the expression. We put up an equation: 660 = 15x + 420, solve for x, and find the value of 30x(adult tickets cost)

Parth (parthkohli):

@lgbasallote Can I have your worthy medal now lol

OpenStudy (lgbasallote):

you can have the asker's more worthy medal :p

OpenStudy (anonymous):

THANKS A MIL PARTHKOLI! Your working and explanation is superb. Had most of it but wanted to be sure!

Parth (parthkohli):

You're welcome @shipwreck \(\Large \color{purple}{\rightarrow x + {1x \over 2} = smartscore }\) \(\Large \color{purple}{\rightarrow {3 \over 2}x }\) @lgbasallote If you gimme the medal, I'll have a profit of 0.5x